Find the slope of the line 2x+4y=5
Answer by lwsshak3(11628)

Find the slope of the line 2x+4y=5
Equation of a line: y=mx+b, m=slope, b=y-intercept
2x+4y=5
4y=-2x+5
y=-2x/4+5/4
y=-x/2+5/4
slope=-1/2